OpenStudy (awy):

Prove the identity. (tanx/1−cosx) = (cscx)(1+secx)

OpenStudy (anonymous):

rewrite the RHS in terms of sin and cosine..

OpenStudy (anonymous):

\[\frac{ 1 }{ sinx}(1+\frac{ 1 }{\cos x})\]

OpenStudy (anonymous):

\[\frac{ 1 }{sinx}(\frac{ \cos x+1 }{ \cos x })\]

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

We can easily convert LHS into RHS...

OpenStudy (anonymous):

Let us start from LHS Rationalize it by multiplying it with (1 + cosx).. \[\frac{\tan(x)}{1 - \cos(x)} \times \frac{1 + \cos(x)}{1 + \cos(x)} = \frac{\tan(x)(1 + \cos(x))}{1 - \cos^2(x)}\]

OpenStudy (anonymous):

\[\implies \frac{\tan(x)(1 + \cos(x))}{\sin^2(x)} \implies \frac{\sin(x)(1 + \cos(x))}{\cos(x) \cdot \sin^2(x)} \implies \frac{cosec(x)(1 + \cos(x))}{\cos(x)}\]

OpenStudy (anonymous):

Using : \[\frac{1}{\sin(x)} = cosec(x)\] \[\frac{1}{\cos(x)} = \sec(x)\]

OpenStudy (anonymous):

\[\implies {cosec(x)} \times (\sec(x) + \sec(x) \cdot \cos(x)) \implies "Hence \; Proved.."\]

OpenStudy (anonymous):

\[\sec(x) \cdot \cos(x) = 1\]